OpenStudy (anonymous):

help hard math word problem! Every sports team at Long Stretch Middle School must select one captain, two junior captains, and one equipment manager. If the softball team has 20 members, how many possible outcomes of the four selections could there be? i think its either b or d A. 20 + 19 + 18 + 17 B. 20 + 20 + 20 + 20 C. 20 × 19 × 18 × 17 D. 20 × 20 × 20 × 20

OpenStudy (dumbcow):

none of them are right lets say there are 20 ways of choosing the captain then that leaves 19 ways of choosing equipment manager then with remaining 18 you must choose 2 junior captains...18 choose 2 \[\rightarrow 20*19*\frac{18!}{2!*16!} = 20*19*17*9\]

OpenStudy (dumbcow):

i would go with C though, which is assuming that the order you pick the junior captains matter